OpenStudy (anonymous):

a ball rolls 2.7 m in 0.55 sec how do you find how far it will rol in 1 sec?

OpenStudy (anonymous):

v = d/t v = 2.7m / 0.55s = 4.9 m/s that means it rolss 4.9m every second

OpenStudy (anonymous):

velocity=displacement/ time

OpenStudy (anonymous):

is there something you don't understand in this problem?

OpenStudy (anonymous):

displacement= the distance which an object has moved

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hiya kush:)

OpenStudy (anonymous):

I suspect Lammy has fallen asleep with this page open on their browser :)

OpenStudy (anonymous):

:) lol

OpenStudy (anonymous):

U frm india aditya?

OpenStudy (anonymous):

yeah. I'm from Cricket world cup 2011 winners :D

OpenStudy (anonymous):

:D ...Me too:)

OpenStudy (anonymous):

High five!

OpenStudy (anonymous):

High five aditya :) !!!

OpenStudy (anonymous):

So , r u a student?

OpenStudy (anonymous):

No, I finished post graduation last year. I am here to learn algorithms and I jump in to the math forum now and then to help out :)

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Hello kush!

OpenStudy (anonymous):

okay,...@aditya yes kush what's the question ???

OpenStudy (anonymous):

HELLO